### v1

# ComGPT: Detecting Local Community Structure with Large Language Models

Abstract
Large language models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks such as node classification. So far, how to leverage LLMs to better detect local communities remains underexplored. Local community detection algorithms based on seed expansion often face a seed-dependent problem, community diffusion, and free rider effect. Using LLMs to solve existing local community work problems faces the following challenges: existing graph encoding methods fail to provide LLMs with sufficient community-related graph information; LLMs lack domain knowledge in mining communities. To address these issues, we improve graph encoding by supplementing community knowledge to enhance the ability of graph encoding to express graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics, aiming to alleviate the seed-dependent problem, community diffusion, and free rider effect. Based on the graph encoding and NSG prompt, we present a GPT-guided local community detection, called ComGPT. ComGPT iteratively selects potential nodes from the detected community’s neighbors and subsequently employs GPT to choose the node that optimally integrates into the detected community from these selected potential nodes. Experimental results demonstrate that ComGPT outperforms the comparison algorithms, thereby confirming the effectiveness of the designed graph encoding and prompt.
Introduction
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention for its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), $k$-core decomposition (Cui et al. 2014), $k$-truss identification (Faysal et al. 2023), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithm has garnered widespread acclaim for its versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the highest quality function value at each step. However, it often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address these problems, researchers developed different strategies such as identifying core node sets and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018).
In this study, we explore an alternative approach inspired by the the success of large language models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022) such as theory-of-mind reasoning (Adhikari et al. 2020; Ammanabrolu and Riedl 2021) and graphical reasoning (Wang et al. 2023). As far as we know, LLMs have not been applied to local community detection. Therefore, we explore the application of large models for local community detection and propose an algorithm, called ComGPT (GPT-guided Local Community Detection). This algorithm leverages GPT to guide the detection process, thereby bridging the current gap in LLMs applicability to this field. ComGPT integrates the seed expansion algorithm with GPT, leveraging their inherent compatibility: GPT through step-by-step guidance, while the seed expansion method incrementally adds nodes to the community, making their integration highly synergistic.
However, the application of LLMs to local community detection faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we enhance graph encoding by incorporating two key aspects: graph topology and community information. 2) Lacks domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we have designed prompt to assist LLMs in understanding community definitions and how to select nodes. 3) Lacks efficient interaction between GPT and community detection methods. The mode for integrating community detection algorithms with GPT, as well as strategies for reducing the number of input tokens during their interaction, remain unexplored. Based on the step-by-step characteristics of GPT and seed expansion algorithm, we design a step-by-step interactive mode. To minimize costs by reducing the number of tokens, the input to GPT is not the entire graph but a partial graph that includes the detected community and some of its neighbors. Our contribution is summarized as follows:
  • We propose a GPT-guided local community detection algorithm, called ComGPT. ComGPT repeatedly selects potential nodes from the neighbors of the detected community and then uses GPT to identify the node that optimally integrates into the detected community from these selected potential nodes. To the best of our knowledge, this is the first work to apply LLMs to the domain of local community detection.
  • For GPT to understand and detect communities, we design graph encoding method, NSG prompt, and interaction mode. The graph encoding method incorporates both graph topology and community information, which enhances GPT to understand the current state of the detected community. NSG prompt that integrates community features facilitates GPT’s comprehension of domain knowledge. Additionally, an interactive mode is established between GPT and the local community detection method.
  • The results on real datasets indicate that ComGPT outperforms comparison methods. Ablation study confirm that GPT plays a crucial role in the ComGPT ’s performance. Furthermore, prompt comparisons experiments reveal that the designed prompt effectively aids GPT in understanding the graph structure and community situation.
Related work
Local community detection
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods like seed expansion algorithms are developed for detecting local communities (Luo, Wang, and Promislow 2008; Cui et al. 2014), which expand communities by selecting nodes with the highest quality function values. For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the highest local modularity R and M, respectively. The seed expansion algorithms often faces the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing GPT, distinguishing it from existing studies.
LLMs
In recent years, researchers have explored LLMs for node classification (Chen et al. 2023) and graph reasoning (Zhang 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect GPT’s interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method tailored for the community detection domain, which incorporates community structure considerations, distinguishing it from existing approaches.
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general COT prompt (Wei et al. 2022) and the BAG prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above work.
Proposed algorithm
Shortcomings
Existing heuristic algorithms suffer from the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013), which are explained as following.
(a) Seed-dependent problem
(b) Community diffusion
(c) Free rider effect
Figure 1: Existing shortcomings. Nodes in the same community are marked with the same color.
Shortcoming 1: Seed-dependent problem (Kloumann and Kleinberg 2014).
It refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(1(a)). The communities detected by the M method (Luo, Wang, and Promislow 2008) with nodes 5 and 1 as the starting seed nodes are [5, 11, 6, 7, 17] and [1, 12, 13, 4, 8, 2, 14, 3, 22, 18, 20], respectively. The former includes some nodes of the real community and misses many nodes, while the latter is more accurate.
Shortcoming 2: Community diffusion.
The community diffusion problem means that the detected community contains nodes from different communities. Take node 1 in Figure 1(1(b)) as an example to illustrate the community diffusion problem. The community detected by the M method with node 1 as the starting seed node is [1, 2, 3, 4, 7, 8, 9, 13, 14, 15, 16, 17, 18, 20], which includes nodes from two different communities.
Shortcoming 3: Free rider effect (Sozio and Gionis 2010) (Lang and Andersen 2007) (Tsourakakis et al. 2013).
The free rider effect refers to mistakenly adding unrelated communities to enhance the fitness of a detected community. We use the community derived from seed node 7, as shown in Figure 1(1(c)), to illustrate it. The target community consists of nodes 5, 6, 7, 8, and 9. However, to achieve a higher fitness value, the detected community not only contains nodes in the target community [5, 6, 7, 8, 9] but also erroneously includes all nodes in the unrelated community [1, 2, 3, 4, 10, 11].
ComGPT
Figure 2: The diagram of ComGPT.
We propose a local community detection algorithm based on GPT, called ComGPT. Figure 2 shows the main steps of ComGPT. It includes four main steps: 1) potential nodes identification; 2) graph encoding; 3) node selection by GPT; 4) node supplementation by GPT. Potential nodes identification aims to pinpoint potential nodes from the neighbors of the current community. Graph encoding involves transforming the graph structure into a textual format interpretable by GPT. Node selection is that GPT selects the node that optimally integrates into the community from potential nodes. Node supplement involves GPT selecting which nodes to add to the community when no potential nodes remain.
Algorithm 1 ComGPT
Input: Network $G$, Seed node $v_{s}$
Output: Community $C$
1:  $C\leftarrow{\{v_{s}\}}$, $cands\leftarrow\emptyset$
2:  while True do
3:     $PN\leftarrow$ potential nodes identification with $C$
4:     while $PN\neq\emptyset$ do
5:        $gtxt\leftarrow$ graph encoding with $C$ and $N$
6:        $v\leftarrow$ node selection by GPT from $PN$
7:        add $v$ to $C$
8:     end while
9:     add $C$ to $cands$
10:     $v\leftarrow$ node supplementation by GPT
11:     if $v$ is not null then
12:        add $v$ to $C$
13:     else
14:        break
15:     end if
16:  end while
17:  $C\leftarrow$ community with highest $M$ in $cands$
18:  return  $C$
Algorithm 1 shows the overall process of ComGPT. For convenience, we briefly introduce the notations used in this paper. Let $G$ denote the network, $v_{s}$ denote the seed node, $C$ denote the detected community, $v$ denote the node returned by the node selection or node appending steps, $gtxt$ denote the text of the graph structure obtained through graph encoding, $PN$ denote the set of potential nodes, $cands$ store the community each time $PN$ is empty. Initially, community $C$ consists of only one seed node $v_{s}$ and the $cands$ is empty (line 1). First, ComGPT selects potential nodes $PN$ from neighbors of current community $C$ for GPT (line 3). The local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$ and NSG (Node Select Guide) prompt, GPT selects the node $v$ from potential nodes $PN$ and adds it to the current community $C$ (line 6,7). The above steps are repeated until $PN$ is empty. After the potential nodes $PN$ is empty, the community $C$ is added to the candidate communities $cands$ (line 9). When $PN$ is empty, it does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Instead, it could be that the nodes have relatively high degrees, which marginally reduces the $M$ value. To address this scenario, ComGPT identifies nodes through node supplementation (line 10). If GPT identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 12). Otherwise, ComGPT stops community expansion (line 14). After the above process is completed, ComGPT selects the community with the highest $M$ value from $cands$ as the final community (line 17).
Potential nodes identification
To mitigate the high computational cost associated with processing nodes through GPT, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT for selection. Different fitness functions are designed in the seed expansion algorithm to select nodes, which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ to select potential nodes, which is defined as the ratio of the number of internal edges in community $C$ to the number of external edges of community $C$. Nodes that are satisfied $\Delta M=M_{C\cup\{v\}}-M_{C}>0$ are regarded as potential nodes. During the initial expansion of the community, since nodes with high degrees often struggle to achieve $\Delta M>0$, these nodes are not considered potential nodes $PN$. To alleviate this preference, these nodes are added to potential nodes based on their closeness value to the current community $C$. The closeness value $f_{v}$ between node $v$ and the community $C$ is calculated as
$f_{v}=\frac{|C\cap N(v)|}{|C|},$ (1)
$|C|$ represents the number of nodes in the community and $N(v)$ represents the neighbors of node $v$. The value of $f_{v}$ ranges from 0 to 1. Higher values indicate a greater proportion of nodes connected to those within the community.
The initial potential nodes consist of nodes with $\Delta M>0$ and nodes with high $f_{v}$ values. To further reduce the number of nodes processed by GPT, $k$ nodes are selected from them to form a potential node’s set.
Graph encoding
Figure 3: Graph encoding.
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT. The graph data includes two aspects. 1) Graph topological structure. We adopt the Incident method (Fatemi, Halcrow, and Perozzi 2023) to let GPT understand the graph’s topological structure. The Incident method represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure inputted to GPT is a local network, i.e., the subgraph consisting of the nodes of the current community $C$ and $PN$. For clarity, the middle part of Figure 3 illustrates the specific graph encoding: “G describes a graph among nodes:1, 2, 3, 5, 6, 7. Node 1 is connected to nodes 2,3”. To reduce the number of tokens, some conjunctions that do not affect the understanding of GPT are removed. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT’s understanding of community structure. This involves outlining which nodes are within the community, those that are outside, and detailing their specific external connections. For clarity, the lower part of Figure 3 illustrates graph encoding for community $C$ = [1, 2, 3] and potential nodes $PN$ = [5, 6, 7]: “Nodes in the current community:[1,2,3]. The outside nodes contain:[5,6,7]. Node 5 is connected to nodes in community:[1,3]”.
Node selection by GPT
Node selection refers to choosing the most suitable node for community expansion, based on the prompt and graph data. Below, we introduce node selection through prompts that teach GPT to accurately select nodes.
1. The more a potential node is connected to other potential nodes, the higher the likelihood of its selection: This prompt could alleviate the seed-dependent problem and the free rider effect.
The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(1(a)) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Taking Figure 1(1(a)) as an example, the community includes node 5, with nodes 1, 7, and 11 regarded as potential nodes. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
The reason for the free rider effect is the neglect of the connectivity of neighbor nodes of the detected community. Nodes that connect to many other neighbor nodes are consideration of their association with other potential nodes, which alleviates the free rider effect. Taking Figure 1(1(c)) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
2. Prioritize selecting nodes that are connected to multiple nodes within the community: This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(1(b)) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where node 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
The first prompt is to consider the closeness with neighboring nodes outside the community, and the second prompt is to consider the closeness with nodes within the community. This is consistent with the characteristics of the community being close inside and sparse outside, thus helping GPT understand the characteristics of the nodes to be selected.
Node supplementation by GPT
When $PN$ is empty, ComGPT uses GPT to supplement nodes. Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuation is warranted, it also determines which additional nodes should be selected. In terms of graph encoding, the graph encoding method mentioned above, including graph topological structure and supplementary knowledge, is used. The difference is that to provide GPT with sufficient information to make a decision, the input graph text corresponds to graph topology that includes the current community and its first and second-order neighbors. The nodes to be selected by GPT are the three nodes with the largest M values among the first-order neighbor nodes of the community. In terms of prompt, it is based on the definition of the community and whether these nodes can maintain the cohesion of the community. If GPT determines that any of these three nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns NULL.
Experiments
To evaluate the performance of our algorithm, we conducted some experiments on real datasets. We first outline the experimental settings, which include datasets, comparison algorithms, and evaluation metrics. We then present the experimental results and provide corresponding analyses.
Experimental settings
dataset M R DMF_M DMF_R LCDNN LCDMD ComGPT
Dolphins 0.4685 0.4505 0.7255 0.7242 0.7558 0.7235 0.7885
Polbooks 0.5257 0.5030 0.7295 0.7286 0.7463 0.7301 0.7586
Football 0.7146 0.7107 0.8881 0.8893 0.8948 0.8134 0.8406
Amazon 0.7706 0.7554 0.8353 0.8361 0.8372 - 0.8385
DBLP 0.4185 0.4185 0.4032 0.4009 0.3152 - 0.4247
Table 1: Comparisons with other methods in terms of fscore.
Datasets
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms.
The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes within this network represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges.The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. In this network, nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors, and edges indicate that two authors have co-authored a paper.
Comparison algorithms
To evaluate the performance of ComGPT, we compare it with five local community detection algorithms: M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LCDMD (Guo et al. 2022). ComGPT uses gpt-3.5-turbo-0125 and sets $k_{f}$ to 0.5. Due to monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
Evaluation metrics
The evaluation metrics precision, recall, fscore (Powers 2020) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. fscore is the harmonic mean of precision and recall. The values of precision, recall, and fscore are between 0 and 1. The larger value implies a better algorithm performance.
Experimental results
Results on networks
For the Dolphins, Polbooks, and Football datasets, each node is selected as a seed node separately. For the Amazon and DBLP datasets, We randomly select 200 nodes from the top 5,000 communities. The average fscore of all seed nodes is calculated. The detailed experimental results are shown in Table 1. Due to the unavailability of the LCDMD (Guo et al. 2022) source code, results for the Dolphins, Polbooks, and Football datasets are cited from (Guo et al. 2022), in which results for the Amazon and DBLP datasets are not available.
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LCDMD on most datasets. ComGPT is superior to other methods because GPT has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, LCDNN, and LCDMD perform better than the M method and R method. The reason why DMF_M, DMF_R, LCDNN, and LCDMD outperform the M method and R method on most datasets is that the latter two methods maximize local modularity (M and R), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, LCDNN, and LCDMD implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize dynamic membership functions to select nodes, which tend to favor nodes with large degrees. LCDNN expands the community based on the fuzzy relationships between nodes and NGC nodes. Meanwhile, LCDMD introduces a new modularity density to effectively mine the core parts of a community.
Ablation study
Method GPT SK NSG precision recall fscore
1 NO NO NO 0.6824 0.7246 0.6978
2 YES NO YES 0.7571 0.8277 0.7832
3 YES YES NO 0.5911 0.7026 0.6221
ComGPT YES YES YES 0.8440 0.8520 0.8445
Table 2: Results of ablation experiment on the football dataset.
To explore the necessity of GPT, SK (Supplementary knowledge in graph coding), and NSG prompt in ComGPT, we conduct ablation experiments on the Football dataset with 50 randomly selected seed nodes. Three comparative models are constructed as follows: 1) Method 1 does not use GPT. The first node in $PN$ will be selected for community expansion until $PN$ is empty; 2) Method 2 only maintains GPT and NSG prompt, without SK; and 3) Method 3 only maintains GPT and SK, without NSG prompt. When calculating precision and recall, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. In this calculation, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
Table 2 shows the experimental results of Methods 1, 2, 3, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, and 3, indicating that GPT, SK, and NSG prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of precision, recall, and fscore, which demonstrates the effectiveness and importance of GPT. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3, suggesting that NSG prompt can significantly enhance its performance in local community detection.
Prompt comparisons
 Method Dolphins Polbooks Football
Zero-shot P 0.7195 0.5793 0.5911
R 0.3911 0.2330 0.7026
F 0.4690 0.3261 0.6221
Few-shot P 0.7136 0.5658 0.6592
R 0.3840 0.3358 0.7600
F 0.4556 0.4070 0.6894
COT P 0.7483 0.5924 0.6120
R 0.4397 0.2633 0.7205
F 0.5173 0.3572 0.6512
BAG P 0.7242 0.6005 0.6530
R 0.4704 0.2338 0.7589
F 0.5192 0.3315 0.6870
NSG P 0.9592 0.7788 0.8440
R 0.7356 0.7559 0.8520
F 0.7964 0.7506 0.8445
Table 3: Results of different prompts. P, R, and F are the abbreviations of recall, precision, and fscore, respectively.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), COT (Wei et al. 2022), and BAG (Wang et al. 2023) prompts. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the indicator values are calculated in the same way as in the ablation study.
Table 3 shows that the performance of NSG is better than that of Zero-shot, Few-shot, COT, and BAG prompts. Specifically, on Polbooks and Football datasets, the Few-shot outperforms the Zero-shot. Few-shot is worse than Zero-shot on the Dolphins dataset. This is because each dataset has a different structure. If the provided examples match the community structure characteristics of the dataset, the performance of GPT is improved. Conversely, if they do not match, GPT’s performance is diminished. The reason COT and BAG outperform Zero-shot on most datasets is that COT enhances GPT’s reasoning ability by guiding it to think step-by-step and BAG helps GPT better digest graph information. On Dolphins, Polbooks, and Football datasets, NSG outperforms other prompts. This is because the design of COT and BAG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
Effect of the number of potential nodes on ComGPT
Figure 4: Effect of $k$ on ComGPT.
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 13, which exhibits higher variability due to the randomness of GPT output. To confirm this, we conduct repeated experiments for $k$ = 13 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
Conclusion
We explored the potential of using GPT to assist in local community detection and developed a GPT-guided methodology, termed ComGPT. Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge. Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge. Experimental results demonstrate that ComGPT outperforms the comparison algorithms. Moreover, the designed graph encoding and prompt help GPT understand the domain knowledge of graph topology and community discovery, respectively.