### v2

# ComGPT: Detecting Local Community Structure with Large Language Models

Abstract
Large Language Models (LLMs), like GPT, have demonstrated the ability to understand graph structures and have achieved excellent performance in various graph reasoning tasks, such as node classification. Despite their strong abilities in graph reasoning tasks, they lack specific domain knowledge and have a weaker understanding of community-related graph information, which hinders their capabilities in the community detection task. Moreover, local community detection algorithms based on seed expansion, referred to as seed expansion algorithms, often face the seed-dependent problem, community diffusion, and free rider effect. To use LLMs to overcome the above shortcomings, we explore a GPT-guided seed expansion algorithm named ComGPT. ComGPT iteratively selects potential nodes by local modularity $M$ from the detected community’s neighbors, and subsequently employs LLMs to choose the node to join the detected community from these selected potential nodes. To address the above issues faced by LLMs, we improve graph encoding method, called Incident, by incorporating community knowledge to improve LLMs’s understanding of community-related graph information. Additionally, we design the NSG (Node Selection Guide) prompt to enhance LLMs’ understanding of community characteristics. Experimental results demonstrate that ComGPT outperforms the comparison methods, thereby confirming the effectiveness of the improved graph encoding method and prompts.
Introduction
Identifying community structures from networks has wide applications, such as identifying protein complexes (Yin et al. 2017). Local community detection has gained attention due to its ability to quickly identify the community containing a specific node (Ni et al. 2024a, b). Various methods are designed for detecting local community, including seed expansion algorithms (Luo, Wang, and Promislow 2008), heat diffusion techniques (Kloster and Gleich 2014), and non-negative matrix factorization algorithms (Wang et al. 2011). Among these methods, the seed expansion algorithms have garnered widespread acclaim for their versatility and ease of implementation. The seed expansion algorithms typically enlarge the community by selecting a node with the high scoring function value at each step (Moradi, Olovsson, and Tsigas 2014).
However, seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013), shown in Figure 1. The seed-dependent problem refers to the quality of the community detected being highly dependent on the initial seed nodes, as illustrated by nodes 5 and 1 in Figure 1(a). The community diffusion problem means that the detected community contains nodes from different communities, as illustrated by node 1 in Figure 1(b). The free rider effect refers to mistakenly adding unrelated communities to enhance the scoring function value of a detected community, as illustrated by node 7 in Figure 1(c). To address these problems, researchers developed different strategies such as identifying core node sets (Guo et al. 2022) and designing scoring functions with a bias towards nodes with large degrees (Luo et al. 2018).
((a))
((b))
((c))
Figure 1: Some shortcomings faced by seed expansion algorithms. Nodes in the same community are marked with the same color.
In this study, inspired by the success of Large Language Models (LLMs) in many fields (Gu et al. 2024; Zong and Krishnamachari 2024; Yang et al. 2022)(Adhikari et al. 2020; Ammanabrolu and Riedl 2021)(Wang et al. 2023), we explore an alternative approach based on LLMs, called ComGPT (GPT-guided seed expansion algorithm). ComGPT leverages LLMs to guide the local community detection process to alleviate the seed-dependent problem, community diffusion, and free rider effect. The one-question-one-answer modality of LLMs aligns well with the incremental nature of seed expansion algorithms. It bridges the current gap in LLMs applicability to the community detection domain.
Specifically, we leverage LLMs to select nodes to join the community instead of the scoring function commonly used in seed expansion algorithms. However, the application of LLMs to select nodes faces the following challenges: 1) Lack of graph encoding methods for providing LLMs with sufficient community-related graph information. To address this, we improve the graph encoding method, called Incident (Fatemi, Halcrow, and Perozzi 2023), by incorporating community information. 2) Lack of domain knowledge for LLMs in dealing with local community detection. It limits the capabilities of LLMs. To address this, we design prompts to assist LLMs in understanding community definitions and how to select nodes. Additionally, LLMs receive only a partial network as input, not the entire network, to minimize costs by reducing the number of tokens processed.
Our contributions are summarized as follows:
  • We propose a GPT-guided seed expansion algorithm called ComGPT, which iteratively uses GPT-3.5-turbo to add nodes to the community. It includes four main steps: potential node identification, graph encoding, node selection by GPT-3.5-turbo, and node supplementation by GPT-3.5-turbo. To the best of our knowledge, this is the first work to apply LLMs to the domain of community detection.
  • We design an improved version of the typical graph encoding method, named Incident (Fatemi, Halcrow, and Perozzi 2023), and NSG (Node Selection Guide) prompt. This improved version incorporates both graph topology and community information, which enhances LLMs to understand the current state of the detected community. NSG prompt that integrates community features facilitates LLMs’s comprehension of domain knowledge.
  • The results on real datasets indicate that ComGPT outperforms comparison methods. The ablation study experiments confirm that LLMs play a crucial role in the ComGPT’s performance. Furthermore, prompt comparison experiments reveal that the designed prompt effectively aids LLMs in understanding the graph structure and community situation.
Related work
Local community detection
Local community detection focuses on discovering the community to which the seed node belongs (Ni et al. 2024c, 2023). A variety of methods, like $k$-core decomposition (Cui et al. 2014) and seed expansion algorithms (Clauset 2005), are developed for detecting local communities. Seed expansion algorithms expand communities by selecting nodes with high scoring function values (Moradi, Olovsson, and Tsigas 2014). For example, the R method (Clauset 2005) and M method (Luo, Wang, and Promislow 2008) choose nodes based on the local modularity R and M, respectively. The seed expansion algorithms often face the seed-dependent problem (Kloumann and Kleinberg 2014), community diffusion, and free rider effect (Sozio and Gionis 2010; Lang and Andersen 2007; Tsourakakis et al. 2013). To address the seed-dependent problem, researchers developed different strategies such as identifying core node sets (Guo et al. 2022; Ji, Guo, and Yu 2022), finding the nearest nodes with greater centrality (Luo et al. 2020), alternating strategy of strong fusion and weak fusion (Shang et al. 2023), and designing membership functions with a bias towards nodes with large degrees (Luo et al. 2018). Our work mitigates the aforementioned problems by utilizing LLMs, distinguishing them from existing studies.
LLMs in graph reasoning
In recent years, researchers have explored the application of LLMs in graph reasoning (Zhang 2023; Chen et al. 2023). Wang et al. demonstrated LLMs’ graph reasoning abilities and designed the NLGraph dataset (Wang et al. 2023) to evaluate these capabilities. Furthermore, Guo et al. emphasized the design and improvement of graph encoding techniques, which critically affect LLMs’ interpretation of input data and its subsequent output (Guo et al. 2023). Therefore, Adjacency and Incident (Fatemi, Halcrow, and Perozzi 2023) are designed to encode graph structured data into text for LLMs. Adjacency uses integer node encoding and parenthesis edge encoding. Incident uses integer node encoding and incident edge encoding. Existing graph coding methods primarily represent graph topology, which is often insufficient for specific tasks, such as community detection. Our work introduces a novel graph encoding method based on Incident (Fatemi, Halcrow, and Perozzi 2023), tailored for the community detection domain, that incorporates community structure and distinguishes itself from existing approaches.
Furthermore, various prompts are designed for specific tasks to make LLMs better adapt to these tasks (Sahoo et al. 2024), such as the general CoT (Chain-of-Thought) prompt (Wei et al. 2022) and the BaG (Build a Graph) prompt (Wang et al. 2023) for graph reasoning tasks. However, these prompts do not incorporate the domain knowledge required for local community detection and are unsuitable for detecting community structures within networks. This paper designs a prompt, named NSG, suitable for the field of local community detection, which is different from the above prompts.
Approach
Figure 2: The diagram of ComGPT. It includes four main steps: 1) Potential node identification aims to pinpoint potential nodes from the neighbors of the current community. 2) Graph encoding involves transforming the graph structure into a textual format interpretable by GPT-3.5-turbo. 3) Node selection by GPT-3.5-turbo is that GPT-3.5-turbo selects the node that optimally integrates into the community from potential nodes. 4) Node supplementation by GPT-3.5-turbo involves GPT-3.5-turbo selecting which nodes to add to the community when no potential nodes remain.
We introduce a GPT-guided seed expansion algorithm named ComGPT, based on GPT-3.5-turbo. This section begins with an overview of ComGPT, followed by detailed descriptions of each step of ComGPT.
Overview
Algorithm 1 ComGPT
Input: Network $G$, Seed node $v_{s}$, Prompts NSG and NSU for node selection and supplementation, respectively, Questions Qsel and Qsup for node selection and supplementation, respectively
Output: Community $C$
1:  $C\leftarrow{\{v_{s}\}}$, $cands\leftarrow\emptyset$
2:  while True do
3:     $PN\leftarrow$ potential nodes identification with $C$
4:     while $PN\neq\emptyset$ do
5:        $gtxt\leftarrow$ graph encoding with ($C\cup PN$, $C$, $PN$)
6:        $v\leftarrow$ node selection by GPT-3.5-turbo with ($gtxt$, NSG, Qsel )
7:        add $v$ to $C$
8:     end while
9:     add $C$ to $cands$
10:     $N\leftarrow$ first and second-order neighbors of $C$
11:     $PN\leftarrow$ nodes selected by modularity $M$
12:     $gtxt\leftarrow$ graph encoding with ($C\cup N$, $C$, $PN$)
13:     $v\leftarrow$ node supplementation by GPT-3.5-turbo with ($gtxt$, NSU, Qsup)
14:     if $v$ is not null then
15:        add $v$ to $C$
16:     else
17:        break
18:     end if
19:  end while
20:  $C\leftarrow$ community with highest $M$ in $cands$
21:  return  $C$
As shown in Figure 2, ComGPT contains four main steps. To better understand the interactions among these steps, Algorithm 1 shows the overall process of ComGPT. Initially, community $C$ contains seed node $v_{s}$ and $cands$ is empty (line 1), where $cands$ stores the community whenever the potential node identification step returns an empty set. First, ComGPT selects potential nodes $PN$ with the step of potential node identification (line 3), used for subsequent graph encoding. In order for GPT-3.5-turbo to understand the graph structure and community information, the local network, i.e., the subgraph consisting of the nodes in both current community $C$ and $PN$ as well as edges between these nodes, is converted into text, i.e., $gtxt$ (line 5). Utilizing $gtxt$, NSG (Node Select Guide) prompt, and Qsel, GPT-3.5-turbo selects node $v$ from potential nodes $PN$ and adds it to the current community $C$ (lines 6-7). The above steps are repeated until $PN$ is empty. If $PN$ is empty, the community $C$ is added to the $cands$ (line 9). The fact that $PN$ is empty does not necessarily imply that the neighboring nodes of the community are weakly connected to the community. Therefore, based on $gtxt$, NSU (Node SUpplementatio) prompt, and Qsup, ComGPT continues to identify nodes through node supplementation (lines 10-13). If GPT-3.5-turbo identifies suitable node $v$ to add to the community, then ComGPT adds $v$ to $C$ and continues to expand current community $C$ (line 15). Otherwise, GPT-3.5-turbo returns “null” and ComGPT stops community expansion (line 17). After the above process is completed, ComGPT selects the community with the highest local modularity $M$ calculated by formula (1) from $cands$ as the final community (line 20).
Potential nodes identification
To mitigate the cost associated with processing nodes through GPT-3.5-turbo, ComGPT identifies potential nodes that may belong to the current community and provides these nodes to GPT-3.5-turbo for selection. Different scoring functions are designed in the seed expansion algorithms to select nodes (Moradi, Olovsson, and Tsigas 2014), which can be used to choose potential nodes, such as modularity $M$ (Luo, Wang, and Promislow 2008), $R$ (Clauset 2005), and $Q_{lcd}$ (Shang et al. 2023). ComGPT uses local modularity $M_{C}$ (Luo, Wang, and Promislow 2008) to select potential nodes. $M_{C}$ is calculated as,
$M_{C}=\frac{e_{ic}}{e_{oc}},$ (1)
where $e_{ic}$ represents the number of internal edges in community $C$ and $e_{oc}$ represents the number of external edges of community $C$. For each first-order neighbor node $v$ of community $C$, if node $v$ satisfies $\Delta M=M_{C\cup\{v\}}-M_{C}>0$, then it is considered a potential node. To further reduce the number of nodes processed by GPT-3.5-turbo, $k$ nodes with the largest $M$ are selected from the above potential nodes to form a potential node’s set $PN$.
Graph encoding
Figure 3: An toy example of graph encoding. The middle part illustrates the text corresponding to graph topological structure. The lower part illustrates the text corresponding to supplementary community information.
The purpose of graph encoding is to convert graph structure data and information about the community $C$ into text that is easily understandable by GPT-3.5-turbo. The graph data includes two aspects. 1) Graph topological structure. We adopt Incident (Fatemi, Halcrow, and Perozzi 2023) to let GPT-3.5-turbo understand the graph’s topological structure. Incident represents the graph topology by introducing the first-order neighbors of each node. To reduce computational overhead, the graph structure input to GPT-3.5-turbo is a local network. For the node selection by GPT-3.5-turbo step, the local network comprises nodes in $C\cup PN$ and the edges between them. For the node supplementation by GPT-3.5-turbo step, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. 2) Supplementary knowledge. We describe node information from a community perspective to enhance GPT-3.5-turbo’s understanding of community structure. This step involves outlining which nodes are within the community, those that are outside, and the connections of outside nodes to both within and outside the community. The outside nodes are potential nodes in $PN$ from the step of potential nodes identification. The text corresponding to the topological structure and the text corresponding to the supplementary information are concatenated to form $gtxt$. To better understanding, an example is given in Figure 3.
Node selection by GPT-3.5-turbo
Node selection refers to choosing the most suitable node for community expansion, based on NSG prompt, graph text $gtxt$, and question Qsel. Graph text $gtxt$ is obtained by the step of graph encoding with the local network, $C$, and $PN$, where the local network consists of the nodes in $C\cup PN$ and edges between these nodes. Below, we introduce NSG prompt and question Qsel.
  • Prompt 1: The more an outside node is connected to other outside nodes, the higher the likelihood of its selection. The outside nodes provided to GPT-3.5-turbo are potential nodes from the step of potential nodes identification. This prompt could alleviate the seed-dependent problem and the free rider effect.
    The reason for the seed-dependent problem is low-degree node preference. Take seed node 5 in Figure 1(a) and M method (Luo, Wang, and Promislow 2008) as an example to illustrate it. Nodes 1, 7, and 11 are each connected to the node in the community [5] by an edge, and node 11, having the smallest degree, is selected so that the community has the fewest external edges. Thus, the M method (Luo, Wang, and Promislow 2008) selects node 11 to join the community. Nodes that connect to many other potential nodes are usually high-degree nodes, which help to avoid bias towards low-degree nodes and thereby alleviate the seed-dependent problem. Nodes 1, 7, and 11 are potential nodes for community [5]. According to prompt 1, node 1 will be selected over the mistakenly considered node 11.
    The reason for the free rider effect is adding nodes that are sparsely connected to the community to the community. Nodes that are connected to multiple potential nodes indirectly maintain close connections to the community, which alleviates the free rider effect. Taking Figure 1(c) as an example, the community contains node 7, with nodes 4, 6, and 8 regarded as potential nodes. According to prompt 1, nodes 6 and 8 have a higher probability of being selected. Node 4 is a node that may cause the free rider effect, and will not be selected.
  • Prompt 2: Prioritize selecting outside nodes that are connected to multiple nodes within the community. The outside nodes mentioned in prompts are potential nodes from the step of potential nodes identification. This prompt could alleviate the community diffusion problem. The reason for the community diffusion is the insufficient consideration of the connections between nodes to be added to the community and the nodes within the community. Therefore, the prompt considers the connection between potential nodes and the community, thereby alleviating the community diffusion problem. Taking Figure 1(b) as an example, the community includes nodes 1 and 3, with nodes 2, 5, 7, 13, 15, and 18 regarded as potential nodes, where nodes 13 or 18 may lead to community diffusion. According to prompt 2, nodes 2 and 5 have a higher probability of being selected, rather than mistakenly selecting nodes 13 or 18.
    The first prompt is to consider the closeness of the node to potential nodes outside the community, and the second prompt is to consider the closeness of the node to nodes within the community. This is consistent with the characteristics of the community being close inside, thus helping GPT-3.5-turbo understand the characteristics of the nodes to be selected. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for selecting nodes is designed: “You’re doing local community detection. Based on the graph data and prompt, please select a node that you think is most likely to belong to the current community $C$ for community expansion. Provide a detailed explanation.”. Based on the above prompts, the text $gtxt$ obtained by graph encoding method, and the question Qsel, GPT selects the node $v$ from potential nodes.
Node supplementation by GPT-3.5-turbo
Node supplementation involves evaluating whether to continue the expansion process or terminate it. If continuing, it also determines which additional nodes should be selected. Next, we introduce the information needed for graph encoding, NSU prompt, and question Qsup in turn. For graph encoding, to provide GPT-3.5-turbo with sufficient information to make a decision, the local network consists of the nodes in $C\cup N$ and edges between these nodes, where $N$ includes first and second-order neighbors of community $C$. Besides, ComGPT forcibly adds the $k$ nodes with the highest $M$ values from the first neighbors of community $C$ to $PN$. NSU prompt is based on the definition of the community and whether these nodes can maintain the cohesion of the community. To guide GPT-3.5-turbo to provide relevant answers, the question Qsel for supplementing nodes is designed: “Please analyze whether these nodes should be added to the community $C$. The probability of not adding nodes is higher. But it doesn’t mean you always refuse to add nodes. If you think there is a suitable node, please output its node number.”. Based on the above prompts, the text obtained by graph encoding method, and the question Qsel, if GPT-3.5-turbo determines that any of nodes warrant inclusion in the community, it returns the node’s number; otherwise, it returns “null”.
Experiments
Experimental settings
 Dataset M R DMF_M DMF_R LCDNN LS ComGPT
Dolphins Jaccard 0.3370 0.3128 0.6270 0.6320 0.6438* 0.7735 0.7035
Fscore 0.4673 0.4478 0.7254 0.7241 0.7558* 0.8385 0.7684
Polbooks Jaccard 0.4266 0.3986 0.6449 0.6437 0.6673 0.6919 0.6468*
Fscore 0.5184 0.4960 0.7294 0.7285 0.7463 0.7764 0.7320*
Football Jaccard 0.6117 0.6134 0.8546 0.8579 0.7547 0.2406 0.7571*
Fscore 0.6706 0.6699 0.8881 0.8893 0.8255* 0.3593 0.8103
Amazon Jaccard 0.6961 0.6771 0.7633 0.7686 0.7665* 0.6506 0.7734
Fscore 0.7706 0.7554 0.8353 0.8361* 0.8372 0.7420 0.8421
DBLP Jaccard 0.3239 0.3207* 0.2945 0.2928 0.2208 0.0650 0.3259
Fscore 0.4185 0.4185 0.4032 0.4009 0.3152 0.0850 0.4359
Table 1: Experimental results on five datasets. On each dataset, the highest Fscore (Jaccard) is highlighted in bold, the second highest one is highlighted by underline, and the third highest one is highlighted with a superscript asterisk.
Datasets
Five real-world datasets, including Dolphins (Lusseau 2003), Football (Girvan and Newman 2002), Polbooks (Krebs 2004), Amazon (Yang and Leskovec 2012), and DBLP (Yang and Leskovec 2012) are utilized to test the performance of algorithms. The Dolphins network is based on observations of associations between pairs of dolphins, where edges represent associations between dolphins. It contains 62 nodes and 159 edges. The Football network contains 115 nodes and 613 edges. Nodes represent college football teams, communities correspond to different leagues, and edges correspond to games played between them. The Polbooks network is a co-purchasing network for books. Each node in the dataset represents a book about US politics. An edge between two books indicates that they are often purchased together by customers. It contains 105 nodes and 441 edges. The Amazon network represents product co-purchasing relationships, consisting of 334,863 nodes and 925,872 edges. Nodes representing products and edges indicate which products are frequently co-purchased. The DBLP network is a co-authorship network comprising 317,080 nodes and 1,049,866 edges. Here, nodes represent authors and edges indicate that two authors have co-authored a paper.
Baselines
To evaluate the performance of ComGPT, we compare it with M method (Luo, Wang, and Promislow 2008), R method (Clauset 2005), DMF_M (Luo et al. 2018), DMF_R (Luo et al. 2018), LCDNN (Luo et al. 2020), and LS (Shi et al. 2024).
Implementation Details
ComGPT uses gpt-3.5-turbo-0125. The number of potential nodes $k$ is set to 10. All experiments are conducted under Windows 10 operating system (CPU: 11th Gen Intel(R) Core(TM) i7-11700K @ 3.60GHZ, Memory: 32 GB.). Due to the monetary costs of GPT-3.5-turbo-0125, we execute ComGPT twice for each starting seed node to handle GPT hallucinations.
Evaluation metrics
The evaluation metrics Fscore (Powers 2011) and Jaccard (Jaccard 1912) are used to evaluate the detected communities. For an overlapped seed node, all the communities it belongs to are integrated as a real local community. Fscore is the harmonic mean of precision and recall. The values of Fscore and Jaccard are between 0 and 1. The larger value implies a better algorithm performance.
Experimental results
Results on networks
For the Dolphins, Polbooks, and Football datasets, all node are selected as seed nodes. For the Amazon and DBLP datasets, we randomly select 200 nodes from the top 5,000 communities as seed nodes. The average Fscore and Jaccard of all seed nodes is calculated. The detailed experimental results are shown in Table 1.
Table 1 shows that ComGPT outperforms the M method, R method, DMF_M, DMF_R, LCDNN, and LS on most datasets. Among all comparison methods, ComGPT exhibits the best performance on the Amazon and DBLP datasets and ranks second or third on the other three datasets. Although LS achieves the best performance on the Dolphins and Polbooks datasets, its performance on the remaining three datasets is the lowest among all comparison methods. Overall, ComGPT is superior to LS. ComGPT is superior to other methods because GPT-3.5-turbo has excellent comprehension to find the suitable node based on the NSG prompt. On the Football dataset, ComGPT performs worse than DMF_M, DMF_R, and LCDNN. It is because Football contains several tightly interconnected communities, which leads GPT-3.5-turbo to erroneously identify them as a single community. Except for the DBLP dataset, DMF_M, DMF_R, and LCDNN perform better than the M method and R method. The reason why DMF_M, DMF_R, and LCDNN outperform the M method and R method on most datasets is that the latter two methods maximize local modularity ($M$ and $R$), making them sensitive to the choice of seed nodes, known as seed-dependent problem. In contrast, DMF_M, DMF_R, and LCDNN implement various strategies to alleviate this problem. Specifically, DMF_M, and DMF_R, utilize membership functions to select nodes with large degrees. LCDNN expands the community based on the fuzzy relations between nodes and NGC nodes.
Ablation study
Method GPT SK NSG NSU Fscore Jaccard
1 NO NO NO NO 0.4666 0.3352
2 YES NO YES YES 0.3711 0.2414
3 YES YES NO YES 0.4855 0.3299
4 YES YES YES NO 0.7134 0.6307
ComGPT YES YES YES YES 0.7688 0.7009
Table 2: Results of ablation experiment on the Dolphins dataset.
To explore the necessity of GPT-3.5-turbo, SK (Supplementary knowledge in graph coding), NSG, and NSU prompts in ComGPT, we conduct ablation experiments on the Dolphins dataset with 50 randomly selected seed nodes. Four simplified versions of ComGPT are constructed as follows: 1) Method 1 without GPT-3.5-turbo. The first node in $PN$ is selected for community expansion until $PN$ is empty; 2) Method 2 without SK; 3) Method 3 without NSG prompt; and 4) Method 4 without NSU prompt. When calculating Fscore and Jaccard, it is necessary to obtain the intersection between the detected community, which may include duplicate nodes, and the true community. Here, duplicate nodes in the detected community are regarded as a single node. To prevent cyclic selection of duplicate nodes, community expansion is terminated once the number of duplicate nodes exceeds one-third of the number of nodes in the community.
Table 2 shows the experimental results of Methods 1, 2, 3, 4, and ComGPT. Overall, ComGPT outperforms Methods 1, 2, 3, and 4, indicating that GPT-3.5-turbo, SK, NSG, and NSU prompt play a crucial role in ComGPT, respectively. ComGPT is better than Method 1 in terms of Fscore and Jaccard, which demonstrates the effectiveness and importance of GPT-3.5-turbo. The reason ComGPT outperforms Method 2 is that its graph coding incorporates community knowledge. Compared with graph encoding methods that contain only the graph topology, supplementing community knowledge enables the GPT-3.5-turbo to better understand the community structure and thus make more accurate judgments. ComGPT also outperforms Method 3 and 4, suggesting that NSG and NSU prompts can enhance its performance in local community detection.
Prompt comparisons
Prompts Description
Few-shot It is given a few demonstrations of the task
at inference time as conditioning.
Zero-shot No demonstrations are allowed and only an instruction in natural language is given to the model.
CoT Let’s think step by step.
BaG Let’s construct a graph with the nodes and edges first.
Table 3: Description of the prompts.
 Prompts Dolphins Polbooks Football
Zero-shot Jaccard 0.3299 0.2191 0.5829
Fscore 0.4855 0.3490 0.6774
Few-shot Jaccard 0.3584 0.2703 0.7145
Fscore 0.5128 0.4078 0.7764
CoT Jaccard 0.3250 0.2078 0.6168
Fscore 0.4785 0.3347 0.7196
BaG Jaccard 0.2772 0.1813 0.6050
Fscore 0.4235 0.2988 0.7131
NSG Jaccard 0.7009 0.6624 0.8045
Fscore 0.7688 0.7418 0.8555
Table 4: Results of different prompts.
To test the performance of the NSG prompt within ComGPT, we replaced the NSG prompt with four typical prompt approaches: Zero-shot (Brown et al. 2020), Few-shot (Brown et al. 2020), CoT (Wei et al. 2022), and BaG (Wang et al. 2023) prompts, as shown in Table 3. These alternate versions of the algorithm serve as baseline models for comparison. We conducted comparative experiments by randomly selecting 50 seed nodes in the Polbooks, Dolphins, and Football datasets. The detected communities may contain duplicate nodes, and the metrics values are calculated in the same way as in the ablation study.
Table 4 shows that the performance of ComGPT with NSG is better than that of ComGPT with Zero-shot, Few-shot, CoT, and BaG prompts. Specifically, on Football dataset, the ComGPT with CoT and BaG outperform the ComGPT with Zero-shot. The ComGPT with CoT and BaG is worse than ComGPT with Zero-shot on Dolphins and Polbooks datasets. Advanced prompts may have limited performance improvement or even counterproductive effects on complex graph reasoning tasks. For instance, advanced prompts like CoT increase the number of reasoning steps for GPT-3.5-turbo, which potentially raises the likelihood of errors in the final reasoning conclusion due to mistakes in one of the intermediate steps. The reason ComGPT with Few-shot outperforms ComGPT with Zero-shot on most datasets is that we provided some examples of how the task should be executed. These examples guide the model by showing the correct process needed to solve the task. On Dolphins, Polbooks, and Football datasets, ComGPT with NSG outperforms ComGPT with other prompts. This is because the design of CoT and BaG is oriented towards universal graph reasoning problems so that no relevant guidance or knowledge is utilized in dealing with local community detection. NSG is oriented to the field of local community detection and utilizes the guidance in local community detection.
Effect of the number of potential nodes on ComGPT
Figure 4: Effect of $k$ on ComGPT.
We explore the effect of the number of potential nodes, denoted by $k$, on ComGPT on the Football dataset, with 50 seed nodes selected randomly. The values of $k$ are 5, 7, 10, 13, and 15. Figure 4 shows the performance of ComGPT under different $k$. As the $k$ increases, the performance of ComGPT remains generally stable, though with some fluctuations. Fluctuations are relatively low across different values, except for $k$ = 10, which exhibits higher variability due to the randomness of GPT-3.5-turbo output. To confirm this, we conduct repeated experiments for $k$ = 10 and the results are nearly identical to those obtained with other values of $k$. The above analysis suggests that the performance of ComGPT is less affected by the parameter $k$.
Conclusion
We explored using GPT-3.5-turbo to assist in seed expansion algorithm and developed a GPT-guided methodology, termed ComGPT. Additionally, a graph encoding method suitable for local community detection is explored, which enhances the existing graph encoding by incorporating community knowledge. Furthermore, the NSG prompt is designed to enhance LLMs’ understanding of community domain knowledge. Experimental results confirm that the designed graph encoding method and prompts help GPT-3.5-turbo understand the domain knowledge and detect local community structure, respectively.